OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@BrianaNicole16 you should google it so u know the process and it has to do with multiplication or division not quite sure :P

OpenStudy (lynfran):

hint: they both shd have 2 solutions

OpenStudy (lynfran):

|5x+6|=41 (5x+6)=41 and (5x+6)=-41 5x=41-6 5x=-41-6 x=35/5 x=-47/5 |2x+13|=28 (2x+13)=28 and (2x+13)=-28 2x=28-13 2x=-28-13 x=15/2 x=-41/2

OpenStudy (anonymous):

what answer choice is that?